# Text2AC-Zero: Consistent Synthesis of Animated Characters using 2D Diffusion

### v1

@@ Abstract
We propose a zero-shot approach for consistent Text-to-Animated-Characters synthesis based on pre-trained Text-to-Image (T2I) diffusion models. Existing Text-to-Video (T2V) methods are expensive to train and require large-scale video datasets to produce diverse characters and motions. At the same time, their zero-shot alternatives fail to produce temporally consistent videos. We strive to bridge this gap, and we introduce a zero-shot approach that produces temporally consistent videos of animated characters and requires no training or fine-tuning. We leverage existing text-based motion diffusion models to generate diverse motions that we utilize to guide a T2I model. To achieve temporal consistency, we introduce the Spatial Latent Alignment module that exploits cross-frame dense correspondences that we compute to align the latents of the video frames. Furthermore, we propose Pixel-Wise Guidance to steer the diffusion process in a direction that minimizes visual discrepancies. Our proposed approach generates temporally consistent videos with diverse motions and styles, outperforming existing zero-shot T2V approaches in terms of pixel-wise consistency and user preference.
Prompt: “A mechanical cyborg moves to the left on a sandy beach” [Uncaptioned image] Prompt: “An astronaut jumps on the moon surface” [Uncaptioned image] Prompt: “A man shimmy dances in a studio”
Figure 1: Text2AC-Zero produces temporally consistent videos of animated characters using pre-trained Text-to-Image diffusion.
Figure 2: Latent diffusion models for Text-to-Image (T2I) synthesis are sensitive to changes in the latent codes. For instance, if we condition Text2Video-Zero [13] on a depth map and then shift this map upwards by 10 pixels, the textures on the generated character (the robot) change dramatically. By examining the distributions of the initial latent codes that are aligned with the character, we observe a distributional shift, as shown on the plot to the right. Our proposed approach attempts to align the latent codes in a zero-shot manner, eliminating the distribution shift and producing consistent images, as shown at the bottom right.
@@ 1 Introduction
Human synthesis is an essential task in computer vision and graphics, with several applications in movie production, art, and fashion. It aims to generate high-quality and diverse images/videos of human characters that adhere to some given conditions. For instance, in fashion, several approaches were proposed to perform novel view synthesis conditioned on a human pose [1, 21], virtual try-on given a new style [3], and video generation based on textual prompts [11]. Even though these approaches can produce high-quality and realistic output, they are domain-specific and are limited to specialized datasets [15]. More specifically, they can generate images/videos of fashion performers in controlled settings but not diverse characters in complex environments.
Text-to-Image (T2I) diffusion models [22, 20, 23] revolutionized the task of image synthesis by harnessing large-scale datasets with billions of image/text pairs [24]. Additional conditioning signals such as human pose and depth maps introduced in ControlNet [31] and T2I-Adapters [18] imposed further control over T2I models. This allowed the generation of highly diverse, faithful, and realistic images of human characters based on various conditioning signals. On the other hand, generating videos of human characters using Text-to-Video (T2V) diffusion models still lags behind due to the complexities associated with training these models. These complexities include modeling the temporal dimension, finding sufficiently large datasets, and fulfilling their excessive computational needs. As an example, the largest publicly available video dataset encompasses only 10 million videos [2], and it requires up to 48 A100-80GB GPUs to train the temporal module of VideoLDM [4].
Alternatively, several one and zero-shot approaches [29, 5, 29] leveraged T2I models to generate videos. Nevertheless, the output from these approaches lacks temporal consistency, and the generated contents tend to drift and flicker. To examine this behavior, we conduct a controlled experiment where we generate a short video using Text2Video-Zero [13] conditioned on depth maps. We obtain these depth maps by rendering a SMPL human model [16] and shifting it upwards by 10 pixels to simulate a moving human in a video, Figure 2 illustrates this experiment, and it shows that Text2Video-Zero fails to produce temporally consistent images even with this simple change to the conditioning signal. We find that this is caused by the distributional shift in the latent codes that are responsible for generating the character in the scene as shown on the right of Figure 2. This is similar to the findings of [4, 8], where they trained a temporal module on video datasets to align the latents of a T2I diffusion model. However, as mentioned earlier, training these temporal modules is computationally expensive and requires large-scale video datasets.
@@
In this paper, we propose a zero-shot approach for aligning the latents of T2I diffusion models that allows generating consistent videos of animated characters. We employ text-based human motion diffusion models [28, 6] to generate a sequence of guidance signals given a text prompt. Then, we use a pre-trained T2I model to generate the video frames conditioned on these guidance signals. To achieve temporal consistency between frames, we introduce the Spatial Latent Alignment module that attempts to align the latent codes between video frames based on a pre-computed cross-frame dense correspondence. Finally, we propose a Pixel-Wise Guidance strategy that steers the diffusion process in a direction that minimizes the visual discrepancies between frames. To evaluate temporal consistency, we introduce the Human Mean Squared Error metric that measures the pixel-wise difference of the animated character between consecutive frames. Our proposed approach outperforms existing zero-shot T2V approaches on this metric by $\sim 10\%$ and was preferred by 76% of the users in a user study that we conducted.
@@
  • We introduce a zero-shot video synthesis approach for animated characters that leverages pre-trained motion and T2I diffusion models.
  • We propose the Spatial Latent Alignment and Pixel-Wise Guidance modules that boost temporal consistency between frames based on cross-frame dense correspondences that we compute.
@@ 2 Related Work
Human Video Synthesis Existing approaches for human video generation are generally limited to specific domains and datasets. For instance, several T2V approaches [17, 30, 25] train on the UCF-101 dataset [27] that includes videos of humans performing 101 diverse actions. However, the generated videos based on this dataset are low resolution and lack visual diversity. Another category of approaches focused on generating videos of fashion performers and is trained on fashion datasets [11, 15]. For instance, Text2Performer [11] proposed a decomposed human representation into pose and appearance in the latent space of a variational autoencoder. This representation is used alongside a diffusion-based motion sampler to generate consistent high-resolution videos of fashion performers. Nevertheless, their approach can only generate videos of performers with standardized motions on a simple background. Our approach aims to generate videos of any human-like characters (e.g. fictional characters) performing diverse actions in complex environments.
Text-to-Video Diffusion Models Text-to-Image (T2I) diffusion models [23, 20, 22] excelled in generating highly realistic and diverse images based on textual prompts by harnessing large-scale image datasets [24]. With the lack of similarly large video datasets to train T2V counterparts, a growing direction of research attempts to exploit existing T2I models to generate videos. VideoLDM [4] proposed to transform a pre-trained Stable Diffusion model [22] into a T2V model by introducing a temporal module and a video upsampler that are trained on video data. Similarly, Make-a-video [25] extended DALLE-2 [20] to a T2V model by temporally aligning the decoder and the upsampler on video data. However, these two approaches require excessive GPU resources to train and are limited by the scale of the available video datasets.
Tune-a-Video [29] adopts a one-shot paradigm and finetunes a pre-trained T2I model to generate a video given a single video/text pair. Nevertheless, this approach requires a video as an input in addition to the text prompt making it more suitable for video editing or Video-to-Video tasks. Text2Video-Zero [13] introduced a purely text-based zero-shot T2V approach that injects motion dynamics into the latents of a T2I diffusion model. Their approach exploited the fact that the output of diffusion models varies under any changes to the latent codes to generate variations of the first frame. However, the generated frames from their approach lack any motion continuity or temporal consistency. In contrast, our approach produces temporally consistent videos of human characters with continuous motion.
@@ 3 Method
In this section, we first describe the existing pipeline for zero-shot Text-to-Video (T2V) diffusion models that is adopted by Text2Video-Zero [13] and MasaCtrl [5]. Then, we explain our proposed zero-shot approach for generating temporally consistent and continuous videos of animated characters.
@@ 3.3 Cross-Frame Dense Correspondences
DensePose provides pixel-wise embeddings for the human body. Nonetheless, the DensePose detector is not temporally-aware, meaning that a given body point is not guaranteed to have the same embedding in all frames. Moreover, downsampling the DensePose embeddings to the latent resolution encompasses quantization, which will misalign the embeddings further. To tackle this issue, we set up a dense correspondence problem between each two consecutive frames based on the DensePose embeddings. The embedding for frame $i$ is denoted as ${P_{i}=[L_{i},U_{i},V_{i}]}$, where $P_{i}\in\mathcal{P}$, $L_{i}$ has pixel-wise labels for body parts in the range of $[0,24]$, and $U_{i},V_{i}$ are UV-coordinates in the range of $[0,255]$. For each body part $j$, we define a set of pixels that belong to that part as ${Q_{i}^{j}:=\{q\ |\ L_{i}(q)=j\}}$. We form a feature vector for each $q\in Q_{i}^{j}$ and we arrange them in the rows of matrix $\hat{P}_{i}^{j}$:
@@ 3.4 Spatial Latent Alignment
Text2Video-Zero [13] Ours
Prompt: “A man lifts dumbbells at the gym”
Prompt: “A ballerina performing a dance in a studio”
Prompt: “A cyberpunk robot jumps off a grass pitch”
Prompt: “Woody from Toy Story doing a jumpy dance”
Figure 4: A qualitative comparison between our proposed approach and the baseline Text2Video-Zero [13]. Our approach is able to generate consistent shapes and textures compared to the baseline.
@@ 3.5 Pixel-Wise Guidance
Note that we apply this process on the resolution of $256\times 256$ rather than the full resolution of $512\times 512$ as the latter would be computationally expensive. This is caused by the cubic complexity of the Hungarian algorithm that is used in Section 3.3.
@@ 4 Experiments
We evaluate our approach based on two baselines by incorporating our approach into their pipelines. The first baseline is MasaCtrl [5], which is an image editing method that can be used to generate a sequence of consistent images, and the second is Text2Video-Zero [13], a zero-shot approach for video synthesis. We also include some qualitative results for a trained T2V approach, Gen-2 [7].
@@ 4.1 Implementation Details
For both baselines, we use a pre-trained Stable Diffusion [22] version 1.5 with ControlNet [31] depth control to generate $512\times 512$ video frames. For inference, we employ the DDIM sampler [26] with a linear schedule. We use $T=100$ inference steps for Text2Video-Zero and $T=50$ for MasaCtrl. We empirically choose the guidance factor ${\delta=0.01}$ in Equation 8, $A=[0,39],B=[20,69]$ for Text2Video-Zero, and $A=[0,19],B=[20,39]$ for MasaCtrl in Algorithm 1. For motion synthesis, we use the official implementation of Motion Diffusion Model (MDM) [28] with some modifications for rendering and computing DensePose [9]. We conduct all experiments on a single NVIDIA A100 GPU except for Gen-2 [7], where we use the official online demo. Our code will be made publicly available on the project page  [^1] https://abdo-eldesokey.github.io/text2ac-zero/.
@@ 4.2 Qualitative Results
Zero-Shot Comparison Figure 4 shows a qualitative comparison between depth-conditioned Text2Video-Zero and our proposed approach. The first row shows that for Text2Video-Zero, the pants of the man in the first frame gradually change into shorts in the following frames, and the shape of the muscles changes between frames. On the other hand, our approach produces consistent clothing and muscles. In the second row, the body of the ballerina is distorted in the second frame of Text2Video-Zero, and the sleeves change color in the third frame. Our approach fixes the distorted frame and maintains the color of the sleeves. The third row shows that the circular light on the robot’s body changes location between frames in Text2Video-Zero. In contrast, our approach generated consistent textures. Finally, the last row shows the cowboy produced by Text2Video-Zero has an inconsistent texture on the shirt, the boots, and the belt. Our approach produces temporally consistent textures in all these regions. We provide more qualitative examples and comparisons against the other baseline MasaCtrl in the supplementary materials. Motion Guidance Significance To demonstrate the impact of motion guidance produced by MDM, we compare our approach against Text2Video-Zero with no depth conditioning, which produces video by injecting motion dynamics into the latent codes. Figure 5 shows that Text2Video-Zero produces random variations of the scene that do not adhere to the motion in the prompt. For example, the astronaut in the top row is just floating and not jumping, and Ironman in the second row does not move but changes pose. On the other hand, our approach produces consistent videos that adhere to the motion in the prompt. We believe that our approach imposes more control over the generated videos as the animated skeletons from MDM can be used as a reference for many customizable human body models with different shapes, genders, and characteristics.
@@ 4.3 Quantitative Results
$H_{MSE}\downarrow$
User Preference [%]
MasaCtrl [5] [ICCV23]
88.19
34 %
@@
Text2Video-Zero [13] [ICCV23]
84.87
24 %
@@
Table 1: Quantitative comparison between our proposed approach and two baselines. The error reduction percentage is shown
@@
We generated 10 videos of diverse motions and characters and we computed the proposed metric and performed the user study on them. The videos will be provided in the supplementary material. Table 1 shows that our approach outperforms both baselines in terms of $H_{MSE}$ by $\sim 9-10\%$ which demonstrates that the generated characters are temporally more consistent. For the user study, users are asked to select between two videos; one is produced by the baseline and the other by our approach. Table 1 shows that $76\%$ and $66\%$ of the users preferred the videos generated by our approach over Text2Video-Zero and MasaCtrl baselines, respectively.

### v2

# LatentMan : Generating Consistent Animated Characters using Image Diffusion Models

@@ Abstract
We propose a zero-shot approach for generating consistent videos of animated characters based on Text-to-Image (T2I) diffusion models. Existing Text-to-Video (T2V) methods are expensive to train and require large-scale video datasets to produce diverse characters and motions. At the same time, their zero-shot alternatives fail to produce temporally consistent videos with continuous motion. We strive to bridge this gap, and we introduce LatentMan that leverages existing text-based motion diffusion models to generate diverse continuous motions to guide the T2I model. To boost the temporal consistency, we introduce the Spatial Latent Alignment module that exploits cross-frame dense correspondences that we compute to align the latents of the video frames. Furthermore, we propose Pixel-Wise Guidance to steer the diffusion process in a direction that minimizes visual discrepancies between frames. Our proposed approach outperforms existing zero-shot T2V approaches in generating videos of animated characters in terms of pixel-wise consistency and user preference. Project page https://abdo-eldesokey.github.io/latentman/.
Figure 1: LatentMan produces temporally consistent videos of animated characters using pre-trained Motion and Text-to-Image (T2I) diffusion models given only a textual prompt.
Figure 2: Cross-Frame Attention (CFAttn) is adopted by multiple zero-shot T2V approaches to generate globally consistent video frames. However, when the conditioning signal (the depth map) changes, e.g. shifted up, the fine details (shown in the insets) tend to vary between frames. We find that this is caused by the distributional shift of the initial latent codes that are aligned with the character, as shown on the plot to the right. Our proposed approach attempts to align the latent codes in a zero-shot manner, eliminating the distribution shift and producing consistent images. *CN refers to ControlNet
@@ 1 Introduction
Generating visual assets of human characters is a prominent task in the realm of image and video synthesis, with many applications in movie production, art, and fashion. This task aims to generate high-quality and diverse images/videos of human characters that adhere to some given conditions, e.g. textual prompts and human poses. Text-to-Image (T2I) diffusion models [25, 24, 26] revolutionized this endeavor as they can generate high-quality images of human characters conditioned on user-provided textual prompts. ControlNet [36] allowed further control over the generated images through various conditioning signals such as depth maps, human poses, and edge maps.
For generating videos of human characters, Text-to-Video (T2V) diffusion models [3, 2, 11, 28] are evolving rapidly, but there are several complexities associated with them. For instance, learning the motion dynamics (e.g. the human body), finding sufficiently large datasets, and fulfilling their excessive computational needs. As an example, the largest publicly available video dataset encompasses only 10 million videos [1], and it requires up to 48 A100-80GB GPUs to train VideoLDM [3] on this dataset. Therefore, a growing direction of research attempts to democratize this task by leveraging T2I models to generate videos in few- to zero-shot manner.
One category of approaches [8, 34, 23, 32] adopts a Video-to-Video (V2V) scheme that relies on a reference video to generate a target video with modified contents. However, these approaches require the user to provide the reference video, which can be difficult and inconvenient to find. Alternatively, Text2Video-Zero [32] proposed to generate videos based only on a textual prompt where the motion dynamic is simulated by applying translation vectors to the latent codes of the first frame. The temporal consistency was achieved by converting the self-attention modules of the T2I UNet, which encodes the visual style, to cross-frame attention. This enforces the T2I model to generate video frames that are visually consistent. Nonetheless, the generated videos lack any motion continuity and only show random variations of the same object. Moreover, a closer look at the generated frames shows that the temporal consistency is rather global, and fine details tend to change.
To illustrate this observation, we conduct a controlled experiment where we render a SMPL human model [18] to obtain a depth map of a human. We use this depth map and a textual prompt as conditions for ControlNet to generate a reference image. Then, we shift the depth map upwards by 10 pixels to simulate a moving human in a video. We replace self-attention modules with cross-frame attention as in Text2Video-Zero to enforce the T2I model to generate frames with the same style as the reference frame. Figure 2 shows that cross-frame attention successfully preserves the overall style of the frames. However, the fine details of the robot (shown in the insets) tend to change between frames. We find that this is caused by the distributional shift in the latent codes that are responsible for generating the character in the scene as shown on the right of Figure 2.
@@
To this end, we propose a zero-shot approach for generating consistent videos of animated characters based on T2I diffusion models. To produce continuous motion dynamics, we employ text-based human motion diffusion models [31] to generate a sequence of SMPL models given a text prompt. We render these SMPL models to generate a sequence of depth maps that can be used as conditional inputs for ControlNet. This allows generating videos with realistic and continuous animations, unlike Text2Video-Zero. To boost temporal consistency, we compute cross-frame dense correspondence based on DensePose [9], and we use it to align the latent codes between video frames through our Spatial Latent Alignment module. We also propose an additional Pixel-Wise Guidance strategy that steers the diffusion process in a direction that minimizes the visual discrepancies between frames.
To evaluate the temporal consistency of the generated videos, we introduce the Human Mean Squared Error metric that measures the pixel-wise difference of the animated character between consecutive frames. Our proposed approach outperforms Text2Video-Zero on this metric by $\sim 10\%$ and was preferred by 76% of the users in a user study that we conducted.
@@
  • We introduce a zero-shot approach for generating videos of animated characters.
  • We employ Motion Diffusion Models to generate continuous motion guidance based solely on text.
  • We propose the Spatial Latent Alignment and Pixel-Wise Guidance modules that boost temporal consistency.
@@ 2 Related Work
Human Video Synthesis Existing approaches for human video generation are generally limited to specific domains and datasets. For instance, several T2V approaches [19, 35, 28] train on the UCF-101 dataset [30] that includes videos of humans performing 101 diverse actions. However, the generated videos based on this dataset are low resolution and lack visual diversity. Another category of approaches focused on generating videos of fashion performers and is trained on fashion datasets [13, 17]. For instance, Text2Performer [13] proposed a decomposed human representation into pose and appearance in the latent space of a variational autoencoder. This representation is used alongside a diffusion-based motion sampler to generate consistent high-resolution videos of fashion performers. Nevertheless, their approach can only generate videos of performers with standardized motions on a simple background.
Recently, several approaches [20, 33, 12] proposed diffusion models for Image-to-Video (I2V) to animate a human character given a subject image and a sequence of poses that are provided by the user. Contrarily, we address the Text-to-Video (T2V) problem that aims to produce diverse videos of animated characters based solely on a textual prompt. It is worth mentioning that the concurrent work [4] shares similarities with our work as it attempts to generate consistent videos given a sequence of UV maps by aligning the latent codes. But we differ from them in that we only require textual prompts as input and that we follow a different strategy for aligning the latents.
Text-to-Video Diffusion Models Text-to-Image (T2I) diffusion models [26, 24, 25] excelled in generating highly realistic and diverse images based on textual prompts by harnessing large-scale image datasets [27]. With the lack of similarly large video datasets to train T2V counterparts, a growing direction of research attempts to exploit existing T2I models to generate videos. VideoLDM [3] proposed to transform a pre-trained Stable Diffusion model [25] into a T2V model by introducing a temporal module and a video upsampler that are trained on video data. Similarly, Make-a-video [28] extended DALLE-2 [24] to a T2V model by temporally aligning the decoder and the upsampler on video data. However, these two approaches require excessive GPU resources and large-scale datasets to train.
Tune-a-Video [32] adopts a one-shot paradigm and finetunes a pre-trained T2I model to generate a video given a single video/text pair. Nevertheless, this approach requires a video as an input in addition to the text prompt, making it more suitable for video editing or Video-to-Video tasks. Text2Video-Zero [15] introduced a purely text-based zero-shot T2V approach that injects motion dynamics into the latents of a T2I diffusion model. Their approach exploited the fact that the output of diffusion models varies under any changes to the latent codes to generate variations of the first frame. However, the generated frames from their approach lack any motion continuity or temporal consistency. In contrast, our approach employs Motion Diffusion Models [31, 6] to generate continuous motion guidance and introduces two strategies for boosting temporal consistency, especially at fine details.
@@ 3 Method
In this section, we first describe the existing pipeline for zero-shot Text-to-Video (T2V) diffusion models that is adopted by Text2Video-Zero [15] and MasaCtrl [5]. Then, we explain our proposed approach for generating temporally consistent videos of animated characters.
@@ 3.3 Cross-Frame Dense Correspondences
Ideally, the UV maps for the SMPL models or DensePose can be used for this purpose. However, since they need to be downsampled to the resolution of the latent code, the correspondences are lost, and they need to be re-computed. To tackle this issue, we set up a dense correspondence problem between each two consecutive frames based on the DensePose embeddings. We opt for DensePose rather than the UV maps as the former divides the human body into parts, making the correspondence problem cheaper to solve. We denote the DensePose embedding for frame $i$ as ${P_{i}=[L_{i},U_{i},V_{i}]}$, where $P_{i}\in\mathcal{P}$, $L_{i}$ has pixel-wise labels for body parts in the range of $[0,24]$, and $U_{i},V_{i}$ are UV-coordinates in the range of $[0,255]$. For each body part $j$, we define a set of pixels that belong to that part as ${Q_{i}^{j}:=\{q\ |\ L_{i}(q)=j\}}$. We form a feature vector for each $q\in Q_{i}^{j}$ and we arrange them in the rows of matrix $\hat{P}_{i}^{j}$:
@@ 3.4 Spatial Latent Alignment
Figure 4: A qualitative comparison between our proposed approach and the baseline Text2Video-Zero [15]. Our approach is able to generate consistent shapes and textures compared to the baseline. The reference frame is the first frame of the video that defines the appearance of the character.
@@ 3.5 Pixel-Wise Guidance
Note that we apply this process on the resolution of $256\times 256$ rather than the full resolution of $512\times 512$ as the latter would be computationally expensive using the Hungarian algorithm with cubic complexity.
@@ 4 Experiments
We evaluate our approach based on two baselines that adopt cross-frame attention. The first baseline is MasaCtrl [5], which is an image editing method that can be used to generate a sequence of consistent images, and the second is Text2Video-Zero [15], a zero-shot approach for video synthesis. Note that Text2Video-Zero has two variations: a condition-free version and a conditional one based on ControlNet. We compare mainly against the latter, but we provide some examples for the former as well.
@@ 4.1 Implementation Details
For both baselines, we use a pre-trained Stable Diffusion [25] version 1.5 with ControlNet [36] depth control to generate $512\times 512$ video frames. For inference, we employ the DDIM sampler [29] with a linear schedule. We use $T=100$ inference steps for Text2Video-Zero and $T=50$ for MasaCtrl. We empirically choose the guidance factor ${\delta=0.01}$ in Equation 8, $A=[0,39],B=[20,69]$ for Text2Video-Zero, and $A=[0,19],B=[20,39]$ for MasaCtrl in Algorithm 1. For motion synthesis, we use the official implementation of Motion Diffusion Model (MDM) [31] with some modifications for rendering and computing DensePose [9]. We conduct all experiments on a single NVIDIA A100 GPU except for Gen-2 [7], where we use the official demo. The code is publicly available  [^1] https://github.com/abdo-eldesokey/latentman.
@@ 4.2 Qualitative Results
Zero-Shot Comparison Figure 4 shows a qualitative comparison between depth-conditioned Text2Video-Zero and our proposed approach. The figure shows that cross-frame attention adopted by Text2Video-Zero is not sufficient to preserve the fine details of the generated characters. As the conditional depth map changes, the object gets distorted (e.g. the robot torso and legs in the first row), or the texture changes (e.g. the pants in the second row become shorts). On the other hand, our approach successfully maintains the fine details of the generated characters across all frames.
Motion Guidance Significance To demonstrate the impact of motion guidance produced by MDM, we compare our approach against Text2Video-Zero with no depth conditioning, which produces video by injecting motion dynamics into the latent codes. Figure 5 shows that Text2Video-Zero produces random variations of the scene that do not adhere to the motion in the prompt. For example, the astronaut in the top row is just floating and not jumping, and Ironman in the second row does not move but changes pose. On the other hand, our approach produces consistent videos that adhere to the motion in the prompt.
@@ 4.3 Quantitative Results
$H_{MSE}\downarrow$ User Preference [%]
MasaCtrl [5] [ICCV23] 88.19 34 %
@@
Text2Video-Zero [15] [ICCV23] 84.87 24 %
@@
Table 1: Quantitative comparison between our proposed approach and two baselines. The error reduction percentage is shown between brackets.
@@
We generated 10 videos of diverse motions and characters and we computed the proposed metric and performed the user study on them. Table 1 shows that our approach outperforms both baselines in terms of $H_{MSE}$ by $\sim 9-10\%$, which demonstrates that the generated characters are temporally more consistent. For the user study, users are asked to select between two videos; one is produced by the baseline and the other by our approach. Table 1 shows that $76\%$ and $66\%$ of the users (based on 23 users) preferred the videos generated by our approach over Text2Video-Zero and MasaCtrl baselines, respectively.